OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle?

OpenStudy (anonymous):

8==============>V

OpenStudy (anonymous):

first time on here

OpenStudy (freckles):

We are given L and W are the dimensions of the rectangle where L stands for length of the rectangle and W stands for width of the rectangle. We are also given by the first sentence that L is 2+3W or L=2+3W We are also given the the sum of the sides equals 100 meters->can you write this as a mathematical equation?

OpenStudy (anonymous):

i am not that good at algebra

OpenStudy (anonymous):

i can't make that many equations

OpenStudy (freckles):

sum means add can you write an equation expressing the sum of the sides=100 meters

OpenStudy (freckles):

there are 4 sides in a rectangle

OpenStudy (freckles):

opposite sides are equal

OpenStudy (freckles):



OpenStudy (freckles):

can you write the mathematical equation now?

OpenStudy (freckles):

for the sum of the sides is 100 meter?

OpenStudy (anonymous):

L=2+3w/100?

OpenStudy (freckles):

do you know what sum of the sides of the rectangle means?

OpenStudy (freckles):

Do you see the side measurements in the rectangle above that I drew?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

L+W+L+W <---this is suppose to be 100 L+W+L+W is the perimeter of the rectangle (this is just the sum of all the sides) L+W+L+W=100 combine like terms on the left hand side.

OpenStudy (freckles):

then sub in the other equation I gave you

OpenStudy (anonymous):

oh ok. ... ... 2L+2w=100 L=2+3w

OpenStudy (freckles):

yes sub the equation equation you have written there into the first

OpenStudy (anonymous):

... ... 3L+6w=100?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4L+6w=100

OpenStudy (freckles):

in the second equation do you see that L is (2+3w) replace the first L you wrote with (2+3w)

OpenStudy (freckles):